# How Reliable Are Automatic Evaluation Methods for Instruction-Tuned LLMs?

### v2

@@ Abstract
Work on instruction-tuned Large Language Models (LLMs) has used automatic methods based on text overlap and LLM judgments as cost-effective alternatives to human evaluation. In this paper, we perform a meta-evaluation of such methods and assess their reliability across a broad range of tasks. We observe that while automatic evaluation methods can approximate human ratings under specific conditions, their validity is highly context-dependent. Specifically, the simple Rouge-L metric correlates well with human ratings for short-answer English tasks but is unreliable in free-form generation tasks and cross-lingual transfer. The effectiveness of the more advanced method of using GPT-4 as a judge diminishes significantly if reference answers are not included in the prompt, which is the scenario where this method has the potential to provide the most value compared to other metrics. Our findings enhance the understanding of how automatic methods should be applied and interpreted when developing and evaluating instruction-tuned LLMs.
@@ 1 Introduction
The gold standard for the evaluation of generative tasks is human annotations, but these scale poorly due to high costs and time constraints. Consequently, the most common approach for assessing generative LLMs is automated evaluation techniques. Among these, two popular methods are measuring text overlap with Rouge-L Lin (2004) and utilizing existing LLMs as automatic judges Zheng et al. (2023). However, these methods only approximate human judgment, prompting questions about their reliability. While previous research has found that automatic evaluation methods correlate well with human assessments Wang et al. (2022); Zheng et al. (2023), it is important to recognise that these findings generalise over tasks of very different types and in different languages. Additionally, correlation measures may not provide reliable estimates of alignment with human ratings, as they are limited in their ability to deal with ties and constant scores, which are common in human annotations (Deutsch et al., 2023).
In this paper, we provide a thorough analysis of two widely-used automatic methods for approximating human judgments, Rouge-L and LLM-as-a-judge. Additionally, we experiment with BertScore, a semantic text similarity measure, to assess its potential utility. We study the reliability of the three measures across a broad range of English-language tasks. In addition to our studies on English, we perform experiments on Swedish, as an initial study on the reliability of these metrics across languages. Instead of using correlation measures, we employ Pairwise Accuracy (Deutsch et al., 2023) to quantify the alignment with human ratings. Our overall goal is to increase our understanding of the reliability of automatic evaluation methods and to establish guidelines regarding their appropriate usage.
@@
  • We propose using Pairwise Accuracy (PA) with tie calibration Deutsch et al. (2023) to enable robust comparisons between metrics, as we observe a high prevalence of ties ratings which renders common metrics, such as Kendall’s $\tau$ and Spearman’s $\rho$, unreliable.
  • GPT-4 aligns well with human judgments when gold reference answers are available. However, its reliability diminishes in the absence of these references, where it shows an overly positive bias. This is especially problematic for free-form tasks, since GPT-4 is commonly used in such settings.
  • Rouge-L offers a cost-effective alternative to GPT-4 for short-answer tasks, while BertScore shows promising results in long-answer tasks. Our findings indicate that GPT-4, while being the best tool for evaluating generations, can be replaced by faster and far less costly alternatives under certain conditions.
  • We observe a decrease in alignment with humans in non-English tasks for Rouge-L and GPT-4, when it does not have access to gold references, suggesting challenges for lesser-resourced languages.
@@ 2 Related Work
In this section, we will review the research on automatic evaluation of generated text, the increasing use of LLMs as evaluators, and the methods used to assess the alignment of automatic metrics with human preferences.
Automatic Evaluation
@@
When addressing multiple-choice questions, assessing the quality of model outputs appears feasible through standard classification metrics like accuracy and F1-score Li et al. (2023a); Lai et al. (2023). While this can be precise, it is rather strict. Such stringent measures could only provide a fair assessment of the model’s performance if the model does not deviate from the instructed format. It is, however, relatively simple for the model to stray from the required format as the tasks diverge from the training data or get more complex. A more common and more lenient method used in the literature is Rouge-L score Honovich et al. (2023); Wang et al. (2023); Mishra et al. (2022); Yin et al. (2023); Lai et al. (2023); Li et al. (2023b). Rouge-L calculates a score based on the longest common subsequence between the output and the gold answer(s). While this can be a more forgiving measure regarding formatting inconsistencies, it still lacks the sophistication to be effective in tasks where free-from answers are expected.
Evaluation Using LLMs
@@
An increasingly common method for evaluating instruction fine-tuned models is to use powerful LLMs as automatic judges Peng et al. (2023); Gilardi et al. (2023); Chen et al. (2023); Kew et al. (2023). Zheng et al. (2023) propose three different variations: (1) pairwise comparison, which asks the LLM to choose its preferred answer or declare a tie; (2) single answer grading, in which the LLM is asked to assign a score to an answer; and (3) reference-guided grading, in which the model is provided with a reference solution if applicable. An approach similar to the second one is used by M3IT Li et al. (2023b) to evaluate the accuracy, relevance, and naturalness using GPT-4 in a multimodal scenario.
Meta-Evaluation
@@
High costs of human evaluation could explain the relatively limited utilization of manual assessments in the literature, despite its being the gold standard of assessment. The recent work on automating evaluation processes and leveraging LLMs has demonstrated substantial agreement with human ratings. Zheng et al. (2023) show that GPT-4’s judgments align with human evaluations at over 80% agreement, reaching levels comparable to human–human agreement. Zhou et al. (2023) also report agreement levels between GPT-4 and human annotators on a par with human-human agreements. Furthermore, Wang et al. (2022) find a consistently strong correlation between Rouge-L scores with accuracy across different models and task types, indicating Rouge-L is a good proxy for accuracy in classification tasks with short outputs. There are also work conducting meta-evaluation of automatic metrics for chat and summarization Shen et al. (2023); Chiang and Lee (2023) using different criteria. There also has been some work on aligning language model evaluations better with human preferences, such as Liu et al. (2024) and Chan et al. (2024). Furthermore Zhang et al. (2020) show that BertScore correlates better than previous metrics in machine translation, while Hanna and Bojar (2021) try to demonstrate where it fails and where it shines in this context.
@@ 3.1 Data
As our training data, we use the Cleaned Alpaca Dataset [^1] Cleaned Alpaca Dataset, which corrects errors found in the original Alpaca Taori et al. (2023). For testing, we use Natural Instructions v2 (NIv2) Mishra et al. (2022); Wang et al. (2022), which spans a diverse range of tasks, including classification, question answering, free-form text generation, and reasoning. This enables fine-grained testing.
@@
To study the metrics’ reliability across the language dimension, we translate both our training and our test data to Swedish using GPT-3.5-turbo. The prompt template and hyperparameters used for translation can be found in Appendix B. Previous work has shown that the automatic translation of the Alpaca dataset produces high-quality results with low noise levels Holmström and Doostmohammadi (2023); Li et al. (2023a). In addition to the two monolingual train datasets, we create an equally-sized English–Swedish bilingual train set by replacing a random 50% of the samples in the Cleaned Alpaca Dataset with their Swedish translations. Our purpose with this bilingual train set is to conduct a controlled study with a more diverse set of bilingual data.
@@ 3.2 Instruction Tuning
  • LLaMa2 Touvron et al. (2023) is trained on mainly English data; only 0.15% of the pretraining data is Swedish. We use both the 7B and the 13B parameter version of the model.
@@
For instruction tuning, we use the same training settings, hyperparameters, and prompts as Alpaca, and use DeepSpeed Rasley et al. (2020) with the same configuration for all models. For more details regarding the implementation, see Appendix D. The code required to replicate our method is accessible in our GitHub repository [^2] Link will be added in the final version to retain anonymity..
@@ 3.3 Evaluation Methods
Human Gold Standard
To establish the gold standard for our evaluation, we hire three bilingual (English, Swedish) evaluators to assess model outputs based on three criteria: naturalness (how natural and fluent the generated response is), relatedness (whether the response is related to the prompt and follows the required format), and correctness (our main criterion). We ask each annotator to rate each criterion on a Likert scale ranging from 1 (significantly deficient) to 3 (completely proficient). For a detailed description of the annotation process and instructions, see Appendix E. The Kendall’s $\tau$ is 0.74 (averaged over pairs of annotators) and the Fleiss’ $\kappa$ Fleiss and Cohen (1973) is 0.63, indicating a substantial agreement between human annotators.
@@
To compare human ratings to other metrics, we use a variant of the majority vote. More specifically, we compress the three ratings into that score which is assigned by at least two raters, and fall back to a neutral score of 2 in cases where all raters have given different scores. We prefer our method over the obvious alternative of taking an average because it reduces the impact of outlier ratings. For reference, among the human ratings of correctness (our main criterion), there is a majority vote for $93.5\%$ of samples.
@@
Our selection of performance metrics is motivated by the desire to cover commonly used methods on a spectrum from surface-based semantics-based methods. On one end of this spectrum, we use Rouge-L Lin (2004), which measures the textual overlap between a generated response and a reference output as the length of the longest common subsequence. On the other end of the spectrum, we use GPT-4 as a judge Zheng et al. (2023). [^3] Running an evaluation (across all models, tasks, and languages) costs around USD 100 and typically takes 15–20 minutes. Though not significantly cheaper than human evaluations, it certainly surpasses them in terms of time efficiency. Similar to previous work Zhou et al. (2023); Kew et al. (2023), we prompt GPT-4 with the same instructions as given to human evaluators and ask it to rate based on the same criteria on the same Likert scale. We also prompt GPT-4 to provide its reasoning before rating, similarly to Kew et al. (2023), whose framework we use for LLM-as-a-judge evaluations. The prompt template used for evaluations is found in Appendix B. In the standard evaluation setting, the gold labels are included in the prompt as a reference for the model; we mark this setting with the suffix -gold. To ablate the effects of the access to gold labels, we perform additional experiments with these labels excluded; we mark these with the suffix -no-gold. Finally, as a point in-between a purely surface-based and a powerful semantics-based performance metric, we use BertScore Zhang et al. (2020), which quantifies semantic overlap in terms of the cosine similarity between contextual embeddings obtained from pretrained language models.
@@ 3.4 Meta-Evaluation Method
In this study we perform a meta-evaluation of both metrics that produce continuous and ordinal ratings. While Spearman’s $\rho$ and Kendall’s $\tau$ are commonly used for such purposes, these metrics fail to handle tasks with constant score vectors and they do not reward correct predictions of ties. Ties are especially frequent in the human ratings, often following a Likert-scale, which the automatic metrics are benchmarked against.
In response, we have chosen to use Pairwise Accuracy with Tie Calibration ($\mathrm{PA}$) for meta-evaluating metrics. Proposed by Deutsch et al. (2023), $\mathrm{PA}$ addresses the shortcomings of traditional metrics by including mechanisms to explicitly account for the prevalence of ties, thus providing a fairer assessment of metrics.
$\mathrm{PA}$ measures the proportion of correctly ranked pairs, including accurately predicted ties. With values ranging from 0 to 1, the metric is more easily interpreted than traditional correlation metrics such as Spearman’s $\rho$ and Kendall’s $\tau$. $\mathrm{PA}$ includes a tie calibration process by defining a threshold value ($\epsilon$) which specifies what is considered a significant difference between scores. A pair of scores with a difference smaller than $\epsilon$ is considered a tie. This is crucial as some metrics are more likely to produce tied values, as can be seen in Table 1. Tie calibration ensures that comparisons between different metrics are fair, regardless of their inclination to predict ties.
@@
Human Ratings 0.557 $\pm$ 0.162 0.0
GPT-4-gold 0.524 $\pm$ 0.154 0.0
@@
Table 1: The average tie proportion per metric when averaged over tie proportions per task for English tasks.
We study the distribution of ties in our data and observe significant variation for different metrics, as shown in Table 1. The average tie proportion for human ratings is 0.557, serving as our benchmark. In contrast, metrics such as GPT-4, Rouge-L, BertScore exhibit varying degrees of tie proportions. GPT-4 has a similar degree of ties compared to human ratings, while BertScore has considerably lower tie proportions. The significant amount of ties and a constant score vector validates the use of $\mathrm{PA}$ to enable a reliable meta-evaluation of our metrics. As an example of constant vector in our case, in Task 034 (cf. §4), there is a constant score of 1 from human raters, illustrating a scenario that could commonly occur in instruction tuning.
Unlike Deutsch et al. (2023), who calculate $\epsilon$ for each task, we calculate the optimal $\epsilon$ for each metric using data from all tasks. We find that this produces a $\mathrm{PA}$ that better reflects the true correlation between human and metric scores, especially for tasks with only ties or mostly ties. Otherwise, for tasks with only ties, the $\epsilon$ could be as large as the value range for the metric and treat every pair of scores as ties. As shown in Table 1, our metric-level $\epsilon$ correlates well with the amount of ties for the metric. With our pre-calculated $\epsilon$ values, we compute $\mathrm{PA}$ over all models per task, aligning with the No-Grouping setting in Deutsch et al. (2023).

### v4

@@ Abstract
Work on instruction-tuned Large Language Models (LLMs) has used automatic methods based on text overlap and LLM judgments as cost-effective alternatives to human evaluation. In this paper, we perform a meta-evaluation of such methods and assess their reliability across a broad range of tasks. In evaluating how well automatic methods align with human evaluations, correlation metrics are the most commonly employed method despite their inherent limitations when dealing with ties and different scales. To address these shortcomings, we use Pairwise Accuracy as an alternative to standard correlation measures. We observe that while automatic evaluation methods can approximate human ratings under specific conditions, their validity is highly context-dependent. Specifically, the simple Rouge-L metric correlates very well with human ratings for short-answer English tasks but is unreliable in free-form generation tasks and cross-lingual scenarios. The effectiveness of the more advanced method of using GPT-4 as a judge diminishes significantly if reference answers are not included in the prompt, which is the scenario where this method has the potential to provide the most value compared to other metrics. Our findings enhance the understanding of how automatic methods should be applied and interpreted when developing and evaluating instruction-tuned LLMs.
$\dagger$\$$\dagger$\$footnotetext: Equal contribution
@@ 1 Introduction
The gold standard for evaluating generative tasks is human annotation, but this scales poorly due to high costs and time constraints. Consequently, the most common approach for assessing generative LLMs is using automated evaluation techniques. Among these, two popular methods are measuring text overlap with Rouge-L Lin (2004) and utilizing existing LLMs as automatic judges Zheng et al. (2023); however, these methods only approximate human judgment, prompting questions about their reliability. While previous research has found that automatic evaluation methods correlate well with human assessments Wang et al. (2022); Zheng et al. (2023), it is important to recognize that these findings generalize over tasks of very different types and in different languages. Additionally, correlation measures may not provide reliable estimates of alignment with human ratings, as they are limited in their ability to deal with ties and constant scores, which are common in human annotations (Deutsch et al., 2023).
In this paper, we provide a thorough analysis of two widely-used automatic methods for approximating human judgments, Rouge-L and LLM-as-a-judge. Additionally, we experiment with BertScore, a semantic text similarity measure, to assess its potential utility. We study the reliability of the three measures across a broad range of English-language tasks. We also perform experiments on Swedish as an initial study on the reliability of these metrics across languages. Instead of using correlation measures, we employ Pairwise Accuracy (Deutsch et al., 2023) to quantify the alignment with human ratings. Our overall goal is to increase our understanding of the reliability of automatic evaluation methods and to establish guidelines regarding their appropriate usage.
@@
  • We adopt Pairwise Accuracy (PA) with tie calibration Deutsch et al. (2023) to enable robust comparisons between metrics, as we observe a high prevalence of tied ratings which renders common metrics, such as Kendall’s $\tau$ and Spearman’s $\rho$, unreliable.
  • We show that GPT-4 aligns well with human judgments when gold reference answers are available. However, its reliability diminishes in the absence of these references, where it shows an overly positive bias. This is especially problematic for free-form tasks, since GPT-4 is commonly used in such settings.
  • We find that GPT-4, while being the best tool for evaluating generations, can be replaced by faster and far less costly alternatives under certain conditions. In particular, we show that Rouge-L offers a cost-effective alternative to GPT-4 for short-answer tasks, while BertScore shows promising results in long-answer tasks.
  • We observe a decrease in alignment with humans in non-English tasks for Rouge-L and GPT-4 in situations where it does not have access to gold references. This suggests that it could be challenging to use automatic evaluation methods for lesser-resourced languages.
@@ 2 Related Work
We start by reviewing the research on the automatic evaluation of generated text, the use of LLMs as evaluators, and the methods applied to assess the alignment of metrics with human preferences.
2.1 Automatic Evaluation
@@
For short-form tasks such as multiple-choice question answering, assessing the quality of model outputs appears feasible through standard classification metrics like accuracy and F1-score Li et al. (2023a); Lai et al. (2023). While such an evaluation can be precise, it is rather strict and can only provide a fair performance assessment if the model does not deviate from the instructed format. However, this easily happens as the tasks diverge from the training data or get more complex. Surface-level similarity measures such as Rouge-L Honovich et al. (2023); Wang et al. (2023); Mishra et al. (2022); Yin et al. (2023); Lai et al. (2023); Li et al. (2023b) are more forgiving regarding formatting inconsistencies, but still lack the sophistication to be effective in tasks where free-from answers are expected.
2.2 Evaluation Using LLMs
@@
An increasingly common method for evaluating instruction fine-tuned models is to use powerful LLMs as automatic judges Peng et al. (2023); Gilardi et al. (2023); Chen et al. (2024); Kew et al. (2023). Zheng et al. (2023) propose three different variations: (1) pairwise comparison, which asks the LLM to choose its preferred answer or declare a tie; (2) single-answer grading, in which the LLM is asked to assign a score to an answer; and (3) reference-guided grading, in which the model is provided with a reference solution (if available). An approach similar to the second one is used by M3IT Li et al. (2023b) to evaluate the accuracy, relevance, and naturalness using GPT-4 in a multimodal scenario.
2.3 Meta-Evaluation
@@
Human evaluation is the gold standard of assessment in natural language processing, but is not widely used in the literature due to its high costs. Instead, authors have turned to automatic evaluation measures that correlate well with human judgments. Wang et al. (2022) find a consistently strong correlation between Rouge-L scores with accuracy across different models and task types, indicating that it is a good proxy for accuracy in classification tasks with short outputs. For machine translation, Zhang* et al. (2020) show that BertScore is better correlated to human judgments than previous metrics, but Hanna and Bojar (2021) also identify setups where it fails.
The recent work on automating evaluation processes and leveraging LLMs has demonstrated substantial agreement with human ratings. Zheng et al. (2023) show that GPT-4’s judgments align with human evaluations at over 80% agreement, reaching levels comparable to human–human agreement. Zhou et al. (2023) also report agreement levels between GPT-4 and human annotators on a par with human–human agreements. There is also work on the meta-evaluation of automatic metrics for chat and summarization Shen et al. (2023); Chiang and Lee (2023) using different criteria, and on aligning language model evaluations better with human preferences, such as Liu et al. (2024) and Chan et al. (2024).
@@ 3.1 Data
As our training data, we use the Cleaned Alpaca Dataset [^1] https://github.com/gururise/AlpacaDataCleaned, which corrects errors found in the original Alpaca Taori et al. (2023). For testing, we use Natural Instructions v2 (NIv2) Mishra et al. (2022); Wang et al. (2022), which spans a diverse range of tasks, including classification, question answering, free-form text generation, and reasoning. This enables fine-grained testing.
@@
To study the metrics’ reliability across the language dimension, we translate both our training and our test data to Swedish using GPT-3.5-turbo. The prompt template and hyperparameters used for translation can be found in Appendix B. Previous work has shown that the automatic translation of the Alpaca dataset produces high-quality results with low noise levels Holmström and Doostmohammadi (2023); Li et al. (2023a). In addition to the two monolingual train datasets, we create an equally-sized English–Swedish bilingual train set by replacing a random 50% of the samples in the Cleaned Alpaca Dataset with their Swedish translations. Our purpose with this bilingual training set is to conduct a controlled study with more diverse bilingual data.
@@ 3.2 Instruction Tuning
  • LLaMa2 Touvron et al. (2023) is trained on mainly English data; only 0.15% of the pretraining data is Swedish. We use both the 7B and the 13B parameter versions of the model.
@@
For instruction tuning, we use the same training settings, hyperparameters, and prompts as Alpaca, and use DeepSpeed Rasley et al. (2020) with the same configuration for all models. For more details regarding the implementation, see Appendix D.
@@ 3.3 Evaluation Methods
Human Assessment
To establish the gold standard for our evaluation, we hire three bilingual (English and Swedish) evaluators to assess model outputs based on three criteria: naturalness (how natural and fluent the generated response is), relatedness (whether the response is related to the prompt and follows the required format), and correctness (whether the response is correct, which is our main criterion). While there are some tasks for which these criteria may not be applicable (especially correctness), they are well-suited for our chosen set of tasks. We ask each annotator to rate each criterion on a Likert scale ranging from 1 (significantly deficient) to 3 (completely proficient). For a detailed description of the annotation process and instructions, see Appendix E. The Kendall’s $\tau$ is 0.74 (averaged over pairs of annotators) and the Fleiss’ $\kappa$ Fleiss and Cohen (1973) is 0.63, indicating a substantial agreement between human annotators.
@@
To compare human ratings to other metrics, we use a variant of the majority vote. More specifically, we compress the three ratings into that score which is assigned by at least two raters, and fall back to a neutral score of 2 in cases where all raters have given different scores. We prefer our method over the obvious alternative of taking an average because it reduces the impact of outlier ratings. For reference, among the human ratings of correctness (our main criterion), there is a majority vote for $93.5\%$ of samples, providing a robust foundation for our comparisons. We treat the human majority vote as the gold standard and compare other metrics against it.
@@
Our selection of performance metrics is motivated by the desire to cover commonly used methods on a spectrum from surface-based semantics-based methods. On one end of this spectrum, we use Rouge-L Lin (2004), which measures the textual overlap between a generated response and a reference output as the length of the longest common subsequence. On the other end of the spectrum, we use GPT-4 as a judge Zheng et al. (2023). [^2] Running an evaluation (across all models, tasks, and languages) costs around USD 100 and typically takes 15–20 minutes. Though not significantly cheaper than human evaluations, it certainly surpasses it in terms of time efficiency. Similar to previous work Zhou et al. (2023); Kew et al. (2023), we prompt GPT-4 with the same instructions that we give to human evaluators and ask it to rate based on the same criteria on the same Likert scale. We also prompt GPT-4 to provide its reasoning before rating, similarly to Kew et al. (2023), whose framework we use for LLM-as-a-judge evaluations. The prompt template used for evaluations is found in Appendix B. In the standard evaluation setting, the gold labels are included in the prompt as a reference for the model; we mark this setting with the suffix -gold. To ablate the effects of the access to gold labels, we perform additional experiments with these labels excluded; we mark these with the suffix -no-gold. Finally, as a point in-between a purely surface-based and a powerful semantics-based performance metric, we use BertScore Zhang* et al. (2020), which quantifies semantic overlap in terms of the cosine similarity between contextual embeddings obtained from pretrained language models.
@@ 3.4 Meta-Evaluation Method
In this study, we perform a meta-evaluation of both metrics that produce continuous and ordinal ratings. While Spearman’s $\rho$ and Kendall’s $\tau$ are commonly used for such purposes, these metrics fail to handle tasks with constant score vectors or with different rating scales, and they do not reward correct predictions of ties. Ties are especially frequent in Likert-scale human ratings, which the automatic metrics are benchmarked against.
In response, we have chosen to use Pairwise Accuracy with Tie Calibration (PA) for meta-evaluating metrics. Proposed by Deutsch et al. (2023), PA addresses the shortcomings of traditional metrics by including mechanisms to explicitly account for the prevalence of ties, thus providing a fairer assessment of metrics.
PA measures the proportion of correctly ranked pairs, including accurately predicted ties. With values ranging from 0 to 1, the metric is more easily interpreted than traditional correlation metrics such as Spearman’s $\rho$ and Kendall’s $\tau$. PA includes a tie calibration process by defining a threshold value, $\epsilon$, which specifies what is considered a significant difference between scores. A pair of scores with a difference smaller than $\epsilon$ is considered a tie. This is crucial as some metrics are more likely to produce tied values, as can be seen in Table 1. Tie calibration ensures that comparisons between different metrics are fair, regardless of their inclination to predict ties or having different rating scales.
@@
Human Ratings 0.557 $\pm$ 0.162 0.000
GPT-4-gold 0.524 $\pm$ 0.154 0.000
@@
Table 1: The average tie proportion per metric for English-language tasks.
We study the distribution of ties in our data and observe significant variation for different metrics, as shown in Table 1. The average tie proportion for human ratings is 0.557, serving as our benchmark. In contrast, metrics such as GPT-4, Rouge-L, BertScore exhibit varying degrees of tie proportions. GPT-4 has a similar degree of ties compared to human ratings, while BertScore has considerably lower tie proportions. The significant amount of ties and a constant score vector validate the use of PA to enable a reliable meta-evaluation of our metrics. As an example of a constant vector in our case, in Task 034 (cf. §4), there is a constant score of 1 from human raters, illustrating a scenario that could commonly occur in instruction tuning.
Unlike Deutsch et al. (2023), who calculate $\epsilon$ for each task, we calculate the optimal $\epsilon$ for each metric using data from all tasks. We find that this produces a PA that better reflects the true correlation between human and metric scores, especially for tasks with only ties or mostly ties. Otherwise, for tasks with only ties, the $\epsilon$ could be as large as the value range for the metric and treat every pair of scores as ties. As shown in Table 1, our metric-level $\epsilon$ correlates well with the number of ties for the metric. With our pre-calculated $\epsilon$ values, we compute PA over all models per task, aligning with the No-Grouping setting in Deutsch et al. (2023).